OpenStudy (anonymous):

How many years will it take for a $50 investment to grow to $5000 if it is compounded continuously at a rate of 5%? Round to two decimal places. Do not include units in your answer A(t) = P*e^(r*t)

OpenStudy (anonymous):

e^0.05t = 100/50 = 2 0.05t = ln 2 t = ln 2/0.05 = 13.86 @a.suarez4462

OpenStudy (anonymous):

am i right :/ @a.suarez4462

OpenStudy (anonymous):

i bleave i am @a.suarez4462

OpenStudy (anonymous):

nope it was 92.10

OpenStudy (anonymous):

ok how you get that?????? @a.suarez4462

OpenStudy (anonymous):

i didnt i was doing an exam on my computer and after you get the wrong answer it shows the correct one

OpenStudy (anonymous):

Oh I see??? @a.suarez4462